# Unsupervised learning in echo state networks for input reconstruction

### v3

@@ 2.3.3 Numerical experiments
$d_{t}=\begin{cases}\cos(\pi t/50)&1\leq t<400,\\ \cos(\pi t/100)+\sin(\pi t/25)&400\leq t<800,\\ \cos^{9}(\pi t/50)&800\leq t\leq 1200.\end{cases}$ (7)
We used a 50-dimensional ESN with the element-wise hyperbolic tangent function ($\tanh$) as the activation function $\sigma$. Let $A_{0}\in\mathbb{R}^{50\times 50}$ be a matrix with elements being i.i.d. samples obtained from $\mathcal{N}(0,1)$. We set $A=0.9\cdot\lambda(A_{0})\cdot A_{0}$, where $\lambda(A_{0})$ denotes the spectral radius of $A_{0}$. We initialized the elements of the input layer $B\in\mathbb{R}^{50\times 1}$ using i.i.d. samples obtained from $\mathcal{N}(0,0.02)$. We conducted numerical experiments for 10 different implementations of parameters $A$ and $B$. We applied the RLS Algorithm 1 to obtain the readout layer $(W_{\mathcal{R}})_{t}$ without explicitly using the original input $d_{t}$. Thus, $(W_{\mathcal{R}})_{t}r_{t}$ represents the ESN output for IR at time $t$.
Figure 2 shows the results obtained from Algorithm 1 using IR based on UL. We used an input varying every 400 steps (Figure 2(a)). Driven by the input, the reservoir state evolved (Figure 2(b)). As learning progressed, the ESN output approached the original input on average (Figure 2(c)). Despite the varying input, the readout weights exhibited minimal adjustment after 400 steps on average (Figure 2(d)). By conducting numerical experiments, we verified that IR in ESNs can be achieved without explicitly using the original input, thus validating UL.
Figure 2: Numerical example of Algorithm 1 used for input reconstruction based on unsupervised learning. a. Time series $d_{t}$ of the ESN input. b. Example of the reservoir state time series $r_{t}$ of the ESN. Each element of $r_{t}$ is represented in the plot with a different color. c. Average time series of the ESN output $(W_{\mathcal{R}})_{t}r_{t}$ (solid orange line) and its standard deviation across 10 trials (shaded region). The original input time series $d_{t}$ is plotted for reference (black dashed line). d. Average of the time-series readout layer updates $\|(W_{\mathcal{R}})_{t+1}-(W_{\mathcal{R}})_{t}\|_{F}$ (black solid line) and its standard deviation across 10 trials (shaded region).
3 Application of unsupervised input reconstruction
In this section, we examine the applications of IR in ESNs. In Section 2, we formulated IR in ESNs as UL. Thus, problems depending on IR are reformulated as UL problems, which can be treated independently of the original ESN input.
Specifically, we initially examine the replication problem of deterministic dynamical systems governing the generation of the original input. We will demonstrate that dynamical system replication can be formulated as a UL problem.
Next, we examine the noise filtering problem to enhance the performance of IR in ESNs in the presence of noise in the input. We demonstrate that a dynamical system replicated using UL can be employed for effective noise filtering.
@@ 3.1.1 Problem formulations
Let $(X,f^{\circ})$ be a dynamical system, where $X\in\mathbb{R}^{k}$ and $f^{\circ}$ is a function from $X$ to itself. We refer to $f^{\circ}$ as the true dynamical system. Let $\{x_{t}\}_{t=1}^{T}$ be a finite orbit of the dynamical system $(X,f^{\circ})$, such that $x_{1}\in X$ and $x_{t+1}=f^{\circ}(x_{t})$ for all $t=1,\dots,T-1$. The purpose of dynamical system replication is to construct a dynamical system $(Y,\hat{f})$ from a finite orbit $\{x_{t}\}_{t=1}^{T}$ (Figure 3). We informally define that dynamical system replication is successful if $f^{\circ}$ and $\hat{f}$ are equivalent.
@@ 3.1.2 Remarks
In (Pathak et al., 2017), an algorithm for dynamical system replication was proposed. The algorithm is based on IR in ESN with a finite orbit $\{x_{t}\}_{t=1}^{T}$, which is used as the ESN input. Consider an ESN $r_{t+1}=g(d_{t},r_{t})$, where $d_{t}=x_{t}$. We recall that $W_{\mathcal{D}}$ is a readout layer obtained using SL so that it reconstructs the input: $W_{\mathcal{D}}r_{t}\approx d_{t}=x_{t}$. The output of the algorithm is the dynamical system on $Y=\mathbb{R}^{n_{\text{r}}}$, which is defined as follows:
$\hat{f}_{\mathcal{D}}:r\mapsto g(W_{\mathcal{D}}r,r)=\sigma[(BW_{\mathcal{D}}+A)r].$ (8)
By applying some conditions on the ESN, $f^{\circ}$ and $\hat{f}_{\mathcal{D}}$ are equivalent in the sense of the topological conjugacy of dynamical systems, supporting the correctness of the dynamical system replication algorithm (See (Pathak et al., 2018; Lu et al., 2018; Lu and Bassett, 2020; Hart et al., 2020; Kim et al., 2021; Hart et al., 2021; Hara and Kokubu, 2022; Grigoryeva et al., 2023; Hart, 2024) for more details).
In Section 2, we showed that $W_{\mathcal{D}}$ is reformulated as $W_{\mathcal{R}}$, which was defined in (5). Thus, the output of the dynamical system replication algorithm is reformulated as follows:
$\hat{f}_{\mathcal{R}}:r\mapsto g(W_{\mathcal{R}}r,r)=\sigma((BW_{\mathcal{R}}+A)r)=\sigma(\hat{A}r).$ (9)
The calculation of $W_{\mathcal{R}}$ and $\hat{A}$ does not explicitly use the finite orbit $\{x_{t}\}_{t=1}^{T}$ of the true dynamical system. This indicates that, similar to IR, dynamical system replication is also possible within the agent (Figure 3).
@@ 3.1.3 Numerical experiments
We conducted numerical experiments to demonstrate dynamical system replication based on UL. Consider the Lorenz-63 system defined as
$\begin{cases}\dot{\xi}=10(\eta-\xi),\\ \dot{\eta}=\xi(28-\zeta)-\eta,\\ \dot{\zeta}=\xi\eta-\frac{8}{3}\zeta.\end{cases}$ (10)
Let $\phi_{\tau}:[\xi(t),\eta(t),\zeta(t)]^{\top}\mapsto[\xi(t+\tau),\eta(t+\tau),\zeta(t+\tau)]^{\top}$ be the flow map of the Lorenz system (10). We used $\phi_{0.02}$ as the true dynamical system $f^{\circ}$. We numerically calculated a 7000-step orbit $\{x_{t}:=\phi_{0.02}^{t-1}(x_{1})\}_{t=1}^{7000}$ of the true dynamical system $\phi_{0.02}$ using the 4th-order Runge–Kutta method with the initial state $x_{1}=[1,1,1]^{\top}$ and a time step size of $0.02$. We used $\{x_{t}\}_{t=1}^{5000}$ for training and $\{x_{t}\}_{t=5001}^{7000}$ for testing. Additionally, we used a 500-dimensional ESN with the element-wise tanh function as the activation function $\sigma$. Let $A_{0}\in\mathbb{R}^{500\times 500}$ be a matrix with elements being i.i.d. samples from $\mathcal{N}(0,1)$. We set $A=1.2\cdot\lambda(A_{0})\cdot A_{0}$, where $\lambda(A_{0})$ is the spectral radius of $A_{0}$. We set the elements of the input layer $B\in\mathbb{R}^{500\times 3}$ using i.i.d. samples obtaind from $\mathcal{N}(0,0.02)$. We conducted numerical experiments for 10 different implementations of parameters $A$ and $B$, and numerically calculated $W_{\mathcal{D}}$ and $W_{\mathcal{R}}$ using (3) and (5), respectively. Then, we numerically calculated a 2000-step orbit $\{\hat{r}_{t}\}_{t=1}^{2000}$ of $\hat{f}_{\mathcal{D}}$ with the initial condition $\hat{r}_{1}=r_{5001}$. Each orbit of the 500-dimensional dynamical system $\{\hat{r}_{t}\}_{t=1}^{2000}$ of $\hat{f}_{\mathcal{D}}$ was projected onto a 3-dimensional space by calculating $\{W_{\mathcal{D}}\hat{r}_{t}\}_{t=1}^{2000}$ for visualization.
We performed the same procedure for $\hat{f}_{\mathcal{R}}$. We used the inverse map of $\sigma=\tanh$ and the Moore–Penrose inverse of matrix $B$ to calculate $W_{\mathcal{R}}$ using (5). It should be noted that, in the numerical experiments, we used numpy.arctanh and numpy.linalg.pinv employed in the Python package NumPy. All numerical calculations were performed with float64 precision as per the default setting of NumPy.
Figure 4 shows a typical numerical example of the orbits of the dynamical systems. Figures 4(a) and (b) show the orbits of the true dynamical system used for training and testing, respectively. We observe that the projected orbits of both the replicated dynamical system $\hat{f}_{\mathcal{D}}$ (Figure 4(c)) and $\hat{f}_{\mathcal{R}}$ (Figure 4(d)) were similar to the orbit of the true dynamical system (Figure 4(b)) in terms of the Lorenz attractor shape. However, the terminal states were different among the replicated dynamical systems although the initial states were the same (Figure 4(c) and (d)).
@@
Figure 4: Typical examples of the orbits of the dynamical systems. a. Orbit of the true dynamical system $f^{\circ}$ used during training. b. Orbit of the true dynamical system $f^{\circ}$ used during testing. c. Projected orbit of replicated dynamical system $\hat{f}_{\mathcal{D}}$ obtained by using supervised learning. d. Projected orbit of replicated dynamical system $\hat{f}_{\mathcal{R}}$ obtained by using unsupervised learning.
@@
This is because the replicated dynamical systems $\hat{f}_{\mathcal{D}}$ and $\hat{f}_{\mathcal{R}}$ are different because of the numerical errors between $W_{\mathcal{D}}$ and $W_{\mathcal{R}}$ (Table 1, 1st row). The numerical errors are attributed to the numerical approximations of the inverse tanh function (Table 1, 2nd row) and the Moore–Penrose inverse (Table 1, 3rd and 4th rows).
Table 1: Minimum and maximum numerical errors regarding the equivalence between $W_{\mathcal{D}}$ and $W_{\mathcal{R}}$ for 10 trials
@@ 3.2.1 Problem formulations
Let $d_{t}\in\mathbb{R}^{n_{\text{in}}}$ be the original ESN input. Suppose that the readout layer $W_{\mathcal{R}}^{(1)}$ is trained for IR using (5) with the reservoir states $r_{t}^{(1)}$ driven by input $d_{t}^{(1)}$ being contaminated by white Gaussian noise with zero mean and covariance matrix $\Sigma^{(1)}\in\mathbb{R}^{n_{\text{in}}\times n_{\text{in}}}$. Consider the case where the readout layer $W_{\mathcal{R}}^{(1)}$ is kept fixed as prior knowledge (Figure 5(i)). The purpose of noise filtering is to reconstruct the original input $d_{t}$ from a history of reservoir states $r_{t}^{(2)}$ driven by input $d_{t}^{(2)}$ that are contaminated by white Gaussian noise with zero mean and covariance matrix $\Sigma^{(2)}$ (Figure 5(ii)). We focus on the $\text{tr}(\Sigma^{(1)})\leq\text{tr}(\Sigma^{(2)})$ case.
@@ 3.2.2 Remarks
As shown in Section 3.1, the trained readout layer $W_{\mathcal{R}}^{(1)}$ induces the autonomous dynamical system $\hat{f}_{\mathcal{R}}^{(1)}$ defined in (9). By definition, the dynamical system $\hat{f}_{\mathcal{R}}^{(1)}$ is regarded as the evolution law of the reservoir state virtually driven by input $d_{t}^{(1)}$ with a noise intensity lower than that of input $d_{t}^{(2)}$. Thus, intuitively, we can use $W_{\mathcal{R}}^{(1)}$ as prior knowledge in IR. We will demonstrate that, using Kalman filter theory and algorithms, we can utilize this prior knowledge to filter the noise and refine the IR performance. Readers may refer to (Kalman, 1960; Grewal and Andrews, 2001) for a comprehensive understanding of the Kalman filter theory and algorithms.
We used the extended Kalman filter to obtain the filtered reservoir state $r_{t}^{(2\to 1)}$ from the noisy reservoir state $r_{t}^{(2)}$. Then, the output $W_{\mathcal{R}}^{(1)}r_{t}^{(2\to 1)}$ is close to $W_{\mathcal{R}}^{(1)}r_{t}^{(1)}$, resulting in an improved reconstruction performance of the original input $d_{t}$ than that of $W_{\mathcal{R}}^{(1)}r_{t}^{(2)}$. To apply the extended Kalman filter, we define the following state-space model:
$r_{t+1}^{(1)}$ $:=\hat{f}_{\mathcal{R}}^{(1)}(r_{t}^{(1)})+w_{t},$ (11)
$r_{t}^{(2)}$ $:=r_{t}^{(1)}+v_{t},$ (12)
where $w_{t}$ and $v_{t}$ are assumed to be white Gaussian noise with zero mean and covariance matrices $Q\in\mathbb{R}^{n_{\text{r}}\times n_{\text{r}}}$ and $R\in\mathbb{R}^{n_{\text{r}}\times n_{\text{r}}}$, respectively. The Kalman filter algorithm requires the values of $Q$ and $R$. The covariance matrix $Q$ of $w_{t}$ is empirically calculated using $\hat{Q}=\frac{1}{T}\sum_{t=1}^{T}w_{t}w_{t}^{\top}$. Note that, as $\hat{f}_{\mathcal{R}}^{(1)}$ can be obtained using UL (see Section 3.1), the same applies to $\hat{Q}$ using the reservoir state $r_{t}^{(1)}$ provided in prior learning (Algorithm 2).
@@
  Input: $\mathcal{R}_{1,T}^{(1)}$, $\sigma$, $\sigma^{-1}$, $A$, $B$,
@@
  Get $W_{\mathcal{R}}^{(1)}$ from $\mathcal{R}_{1,T}^{(1)},\sigma^{-1},A,B$ using (5)
  Get $\hat{f}_{\mathcal{R}}^{(1)}$ from $W_{\mathcal{R}}^{(1)},\sigma,A,B$ using (9)
@@
However, the covariance matrix $R$ of $v_{t}$ cannot be calculated in prior learning. This is because the empirical estimate of the covariance matrix $R$ of $v_{t}$ requires $r_{t}^{(2)}$ samples. Thus, we used the adaptive Kalman filter algorithm (Mehra, 1972; Mohamed and Schwarz, 1999; Akhlaghi et al., 2017) to adaptively update an empirical estimation of the covariance matrix $R$ of $v_{t}$ based on the samples obtained from the noisy reservoir state $r_{t}^{(2)}$ provided at each time. We denote the update rate in the estimation of the covariance matrix $R$ as $\alpha_{R}\in[0,1]$. Therefore, we propose Algorithm 3, which performs noise filtering without the explicit use of inputs $d_{t}$, $d_{t}^{(1)}$, and $d_{t}^{(2)}$.
Algorithm 3 IR with adaptive Kalman filter
  Input: $r_{1}^{(2)},\dots,r_{T}^{(2)}$, $W_{\mathcal{R}}^{(1)}$, $\hat{f}_{\mathcal{R}}^{(1)}$, $\hat{Q}$, $\alpha_{R}$
@@
     $\hat{r}_{t/t-1}=\hat{f}_{\mathcal{R}}^{(1)}(\hat{r}_{t-1/t-1})$
     $P_{t/t-1}=F_{t-1}P_{t-1/t-1}F_{t-1}^{\top}+\hat{Q}$,
     where $F_{t-1}=\left(\partial\hat{f}_{\mathcal{R}}^{(1)}/\partial r\right)(\hat{r}_{t-1/t-1})$.
     $K_{t}=P_{t/t-1}(P_{t/t-1}+\hat{R}_{t-1})^{-1}$
     $\hat{r}_{t/t}=\hat{r}_{t/t-1}+K_{t}(r_{t}^{(2)}-\hat{r}_{t/t-1})$
     $P_{t/t}=P_{t/t-1}-K_{t}P_{t/t-1}$
     $\hat{v}_{t}=r_{t}^{(2)}-\hat{r}_{t/t}$
     $\hat{R}_{t}=(1-\alpha_{R})\cdot\hat{R}_{t-1}+\alpha_{R}\cdot(\hat{v}_{t}\hat{v}_{t}^{\top}+P_{t/t})$
     $r_{t}^{(2\to 1)}=\hat{r}_{t/t}$

### v4

@@ 2.3.2 Numerical Experiments
$d_{t}=\begin{cases}\cos(\pi t/50)&1\leq t<400,\\ \cos(\pi t/100)+\sin(\pi t/25)&400\leq t<800,\\ \cos^{9}(\pi t/50)&800\leq t\leq 1200.\end{cases}$
We used a 50-dimensional ESN with the element-wise hyperbolic tangent function ($\tanh$) as the activation function $\sigma$. We initialized the elements of the input layer $A$ using i.i.d. samples obtained from $\mathcal{N}(0,0.02)$. Let $B_{0}\in\mathbb{R}^{50\times 50}$ be a matrix with elements being i.i.d. samples obtained from $\mathcal{N}(0,1)$. We set $B=0.9\cdot\rho(B_{0})\cdot B_{0}$, where $\rho(B_{0})$ denotes the spectral radius of $B_{0}$. We conducted numerical experiments for 10 different implementations of parameters $A$ and $B$. We applied the RLS Algorithm 1 to obtain the readout layer $(W_{\mathcal{R}})_{t}$ without explicitly using the original input $d_{t}$. Thus, $(W_{\mathcal{R}})_{t}r_{t}$ represents the ESN output for IR at time $t$.
Figure 2 shows the results obtained from Algorithm 1 using IR based on UL. We used an input that changed every 400 steps (Figure 2(a)). Driven by the input, the reservoir state evolved (Figure 2(b)). As learning progressed, the ESN output approached the original input on average (Figure 2(c)). Despite variations in the input, the readout weights exhibited only minimal changes beyond 400 steps, as averaged over trials with different ESN parameterizations (Figure 2(d)). By conducting numerical experiments, we verified that IR in ESNs can be achieved without explicitly using the original input, thus validating UL.
Figure 2: Numerical example of Algorithm 1 used for input reconstruction based on unsupervised learning. a. Time series $d_{t}$ of the ESN input. b. Example of the reservoir state time series $r_{t}$ of the ESN. Each element of $r_{t}$ is represented in the plot with a different color. c. Average time series of the ESN output $(W_{\mathcal{R}})_{t}r_{t}$ (solid orange line) and its standard deviation across 10 trials (shaded region). The original input time series $d_{t}$ is plotted for reference (black dashed line). d. Average magnitude of the time-series readout layer updates $\|(W_{\mathcal{R}})_{t+1}-(W_{\mathcal{R}})_{t}\|_{F}$ (black solid line) and its standard deviation across 10 trials (shaded region).
3 Applications of Unsupervised Input Reconstruction
In Section 2, we formulated IR in ESNs as a UL problem. This reformulation enables tasks that rely on IR to be addressed within the UL framework, independently of the original input sequence. In this section, we explore two such applications: dynamical system replication and noise filtering.
We first consider the replication of deterministic dynamical systems that generate the input time series. We show that this task can be reformulated as a UL problem.
Then, we address noise filtering, where the goal is to improve the robustness of IR when the input is corrupted by noise. We demonstrate that the replicated system obtained via UL can be used for effective noise suppression.
@@ 3.1.1 Problem Formulations
We begin by formulating the dynamical system replication task. Let $(X,f^{\circ})$ be a dynamical system, where $X\subseteq\mathbb{R}^{k}$ and $f^{\circ}$ is a function from $X$ to itself. We refer to $f^{\circ}$ as the true dynamical system. Let $\{x_{t}\}_{t=1}^{T}$ be a finite orbit of the dynamical system $(X,f^{\circ})$, such that $x_{1}\in X$ and $x_{t+1}=f^{\circ}(x_{t})$ for all $t=1,\dots,T-1$. The goal of dynamical system replication is to construct a dynamical system $(Y,\hat{f})$ based on a finite orbit $\{x_{t}\}_{t=1}^{T}$ (Figure 3). We informally define that dynamical system replication is successful if $f^{\circ}$ and $\hat{f}$ are equivalent, where the notion of equivalence will be made precise in subsequent sections depending on the learning algorithm.
@@ 3.1.2 Remarks
An algorithm for dynamical system replication was proposed in (Pathak et al.,, 2017). The algorithm is based on supervised learning for IR in ESN with a finite orbit $\{x_{t}\}_{t=1}^{T}$, which is used as the ESN input. Consider an ESN $r_{t+1}=g(d_{t},r_{t})$, where $d_{t}=x_{t}$. We recall that $W_{\mathcal{D}}$ is a readout layer obtained using SL so that it reconstructs the input: $W_{\mathcal{D}}r_{t}\approx d_{t}=x_{t}$. The output of the algorithm is the dynamical system on $Y=\mathbb{R}^{n_{\text{r}}}$, which is defined as follows:
$\hat{f}_{\mathcal{D}}:r\mapsto g\left(W_{\mathcal{D}}r,r\right)=\sigma\left[\left(AW_{\mathcal{D}}+B\right)r\right].$ (3.1)
By applying some conditions on the ESN, $f^{\circ}$ and $\hat{f}_{\mathcal{D}}$ are equivalent in the sense of the topological conjugacy of dynamical systems (Katok and Hasselblatt,, 1995), supporting the correctness of the dynamical system replication algorithm (see,e.g., Pathak et al.,, 2018; Lu et al.,, 2018; Lu and Bassett,, 2020; Hart et al.,, 2020; Kim et al.,, 2021; Hart et al.,, 2021; Hara and Kokubu,, 2022; Grigoryeva et al.,, 2023; Hart,, 2024, for more details).
In Section 2, we showed that $W_{\mathcal{D}}$ can be reformulated as $W_{\mathcal{R}}$, which was defined in Equation 2.5. Thus, the output of the dynamical system replication algorithm is reformulated as follows:
$\hat{f}_{\mathcal{R}}:r\mapsto g(W_{\mathcal{R}}r,r)=\sigma\left[\left(AW_{\mathcal{R}}+B\right)r\right]=\sigma\left(\hat{B}r\right).$ (3.2)
The calculation of $W_{\mathcal{R}}$ and $\hat{B}$ does not explicitly use the finite orbit $\{x_{t}\}_{t=1}^{T}$ of the true dynamical system. This indicates that, similar to IR, dynamical system replication is also possible within the ESN architecture (Figure 3).
We further note that the computation of $\hat{B}$ in Equation 3.2, which minimizes the loss function defined in Equation 2.6, does not require access to the ESN parameters $A$ and $B$, but only the activation function $\sigma$. Since computing $W_{\mathcal{R}}$ does require knowledge of $A$ and $B$, this observation implies that, interestingly, dynamical system replication is a strictly simpler task than input reconstruction in terms of internal knowledge requirements.
@@ 3.1.3 Numerical Experiments
We conducted numerical experiments to demonstrate dynamical system replication based on UL. Consider the Lorenz-63 system defined by
$\begin{cases}\dot{\xi}=10(\eta-\xi),\\ \dot{\eta}=\xi(28-\zeta)-\eta,\\ \dot{\zeta}=\xi\eta-\frac{8}{3}\zeta.\end{cases}$
[…]
We used the inverse map of $\sigma=\tanh$ and the Moore–Penrose inverse of matrix $A$ to calculate $W_{\mathcal{R}}$ using Equation 2.5. It should be noted that, in the numerical experiments, we used numpy.arctanh and numpy.linalg.pinv employed in the Python package NumPy. All numerical calculations were performed with float64 precision as per the default setting of NumPy. Figure 4 shows a typical numerical example of the orbits of the dynamical systems. Figures 4(a) and (b) show the orbits of the true dynamical system used for training and testing, respectively. We observe that the projected orbits of both replicated systems, $\hat{f}_{\mathcal{D}}$ (Figure 4c) and $\hat{f}_{\mathcal{R}}$ (Figure 4d), closely resemble the true orbit (Figure 4b) in terms of the Lorenz attractor’s shape. This observation suggests that the replication was successful in preserving the qualitative dynamics, particularly with respect to topological conjugacy. On the other hand, the terminal states differed across the replicated dynamical systems, even though they shared the same initial state (Figure 4(c) and (d)).
@@
Figure 4: Typical orbits of the dynamical systems. a. Orbit of the true dynamical system $f^{\circ}$ (training). b. Orbit of the true dynamical system $f^{\circ}$ (testing). c. Projected orbit of the replicated system $\hat{f}_{\mathcal{D}}$ (supervised learning). d. Projected orbit of the replicated system $\hat{f}_{\mathcal{R}}$ (unsupervised learning), with the same initial condition as in (c).
@@
The discrepancy between $\hat{f}_{\mathcal{D}}$ and $\hat{f}_{\mathcal{R}}$ is attributed to numerical errors in the calculation of $W_{\mathcal{D}}$ and $W_{\mathcal{R}}$ (Table 1, 1st row). The numerical errors stem from the approximations involved in computing the inverse tanh function (Table 1, second row) and the Moore–Penrose inverse (third and fourth rows).
Table 1: Numerical errors in comparing $W_{\mathcal{D}}$ and $W_{\mathcal{R}}$ across 10 trials
@@ 3.2.1 Problem Formulations
We proceed to formulate the noise filtering task in ESNs. Let $d_{t}\in\mathbb{R}^{n_{\text{in}}}$ be the original ESN input. Suppose that the readout layer $W_{\mathcal{R}}^{(1)}$ is trained for IR using Equation 2.5 with the reservoir states $r_{t}^{(1)}$ driven by input $d_{t}^{(1)}$ being contaminated by white Gaussian noise with zero mean and covariance matrix $\Sigma^{(1)}\in\mathbb{R}^{n_{\text{in}}\times n_{\text{in}}}$. Consider the case where the readout layer $W_{\mathcal{R}}^{(1)}$ is retained as prior knowledge (Figure 5(i)). The purpose of noise filtering is to reconstruct the original input $d_{t}$ from a history of reservoir states $r_{t}^{(2)}$ driven by input $d_{t}^{(2)}$ that are contaminated by white Gaussian noise with zero mean and covariance matrix $\Sigma^{(2)}$ (Figure 5(ii)). We focus on the $\text{tr}(\Sigma^{(1)})\leq\text{tr}(\Sigma^{(2)})$ case.
@@ 3.2.2 Remarks
As shown in Section 3.1, the trained readout layer $W_{\mathcal{R}}^{(1)}$ induces the autonomous dynamical system $\hat{f}_{\mathcal{R}}^{(1)}$ defined in Equation 3.2. By definition, the dynamical system $\hat{f}_{\mathcal{R}}^{(1)}$ is regarded as the evolution law of the reservoir state virtually driven by input $d_{t}^{(1)}$ with a noise intensity lower than that of input $d_{t}^{(2)}$. Thus, intuitively, we can use $W_{\mathcal{R}}^{(1)}$ as prior knowledge in IR. We demonstrate that prior knowledge can be effectively leveraged via Kalman filtering to enhance the performance of IR. For a comprehensive understanding of Kalman filter theory and algorithms, we refer the reader to (Kalman,, 1960; Grewal and Andrews,, 2001).
[…]
Then, the output $W_{\mathcal{R}}^{(1)}r_{t}^{(2\to 1)}$ is close to $W_{\mathcal{R}}^{(1)}r_{t}^{(1)}$, thereby yielding a more accurate reconstruction of the original input $d_{t}$ than that obtained from $W_{\mathcal{R}}^{(1)}r_{t}^{(2)}$. To apply EnKF, we define the following state-space model:
$\begin{cases}r_{t+1}^{(1)}&:=\hat{f}_{\mathcal{R}}^{(1)}(r_{t}^{(1)})+w_{t},\\ r_{t}^{(2)}&:=r_{t}^{(1)}+v_{t}.\end{cases},$ (3.3)
where $w_{t}$ and $v_{t}$ are assumed to be white Gaussian noise with zero mean and covariance matrices $Q\in\mathbb{R}^{n_{\text{r}}\times n_{\text{r}}}$ and $R\in\mathbb{R}^{n_{\text{r}}\times n_{\text{r}}}$, respectively. The Kalman filter requires the specification of the covariance matrices $Q$ and $R$. The covariance matrix $Q$ of $w_{t}$ is empirically calculated using $\hat{Q}=\frac{1}{T}\sum_{t=1}^{T}w_{t}w_{t}^{\top}$. Note that, as $\hat{f}_{\mathcal{R}}^{(1)}$ can be obtained using UL (see Section 3.1), the same applies to $\hat{Q}$ using the reservoir state $r_{t}^{(1)}$ provided in prior learning (Algorithm 2).
@@
Input: $\mathcal{R}_{1,T}^{(1)}$, $A$, $B$, $\sigma$
@@
 Get $W_{\mathcal{R}}^{(1)}$ from $\mathcal{R}_{1,T}^{(1)},A,B,\sigma$ using Equation 2.5
 Get $\hat{f}_{\mathcal{R}}^{(1)}$ from $W_{\mathcal{R}}^{(1)},A,B,\sigma$ using Equation 3.2
@@
However, the covariance matrix $R$ of $v_{t}$ cannot be calculated in prior learning. This is because the empirical estimation of $R$ depends on observations of $r_{t}^{(2)}$. Thus, we used the adaptive updating methods (Mehra,, 1972; Mohamed and Schwarz,, 1999; Akhlaghi et al.,, 2017) to calculate an empirical estimation of the covariance matrix $R$ of $v_{t}$ based on the samples obtained from the noisy reservoir state $r_{t}^{(2)}$ provided at each time. We denote the update rate in the estimation of the covariance matrix $R$ as $\alpha_{R}\in[0,1]$. Therefore, we propose Algorithm 3, which performs noise filtering without the explicit use of inputs $d_{t}$, $d_{t}^{(1)}$, and $d_{t}^{(2)}$.
Algorithm 3 IR with adaptive EnKF
Input: $r_{1}^{(2)},\dots,r_{T}^{(2)}$, $W_{\mathcal{R}}^{(1)}$, $\hat{f}_{\mathcal{R}}^{(1)}$, $\hat{Q}$, $M$, $\alpha_{R}$
@@
[…]
  $K_{t}\leftarrow U_{t/t-1}V_{t/t-1}^{-1}$
[…]
  $v_{t}\leftarrow r_{t}^{(2)}-r_{t}^{(2\to 1)}$
  $R_{t}\leftarrow(1-\alpha_{R})\cdot R_{t-1}+\alpha_{R}\cdot(v_{t}v_{t}^{\top}+P_{t/t})$